OpenStudy (anonymous):

If an organism's diploid number is 12, what would be its haploid number? A) 3 B) 6 C) 24 D) 48

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

diploid =2n and haploid=n. If 2n=12 then n= ?

OpenStudy (abhisar):

n=12/2= ?

OpenStudy (anonymous):

C?

OpenStudy (abhisar):

\(\huge\sf \frac{12}{2}=?\)

OpenStudy (anonymous):

oh okay for some reason I just saw your second post "n=12/2= ?" that was weird but its B yeah?

OpenStudy (abhisar):

\(\huge\checkmark\)